# Soft-Clamped Perimeter Modes of Polygon Resonators

### v1

@@ Abstract
Polygon resonators are known for their small size and high force sensitivity. We propose an analytical model to predict the mechanical quality factors ($Q$) of polygon perimeter modes by incorporating a stress term in the Timoshenko-Gere equation. We discuss the origin of losses and their relationship with polygon structure dimensions. Notably, we demonstrate that the torsion loss can be suppressed through the torsion angle of supporting tethers. These results offer new insights into optimizing high $Q$ polygon resonators for cavity optomechanics, force sensing, and multimode physics of phononic structures.
@@
Recent advancements have facilitated ultra-high quality factor ($Q$) nanomechanical resonators, attributed to the phenomenon known as ”dissipation dilution” Tsaturyan (2017); González (2000). This phenomenon stems from the nonlinear stress-strain coupling, which gives rise to a quadratic relationship between the resonator’s elongation and its vibration amplitude Fedorov et al. (2019). Notably, dissipation dilution significantly minimizes energy loss in the mechanical system through geometric engineering and stress enhancement, rather than relying solely on material optimizing Tsaturyan (2017); Fedorov et al. (2019). This opens up exciting avenues for device design.
@@
Investigating the structural design and forces inherent in nanomechanical systems has revealed that the mode curvature at boundaries is a major source of the mechanical energy loss Yu et al. (2012); Tsaturyan (2017). Reducing the amplitude of the mechanical modes is an effective way to minimize the energy loss in these systems. Several innovative techniques, referred to as ”soft clamping” Tsaturyan (2017), have been developed specifically for the clamping points to achieve the low-energy-loss. Previous studies have concentrated on tapered clamping Bereyhi et al. (2019), hierarchical structuring Bereyhi et al. (2022a); Fedorov et al. (2020), phononic bandgap engineering Cupertino et al. (2024); Beccari et al. (2022); Ghadimi (2018), and perimeter modes in polygon resonators Bereyhi et al. (2022b).
@@
Remarkably, the perimeter modes in polygon resonators have emerged as promising candidates for integrated quantum optomechanical systems Bereyhi et al. (2022b). In the tapered-clamping approach Bereyhi et al. (2019), enhancement of $Q$ is challenging, as it is constrained to the material’s yield stress-to-deposition stress ratio. In the hierarchical structuring approach Bereyhi et al. (2022a); Fedorov et al. (2020), the torsional losses in cascaded branchings hinder the resonators from attaining the ”clampless” limit. Furthermore, in comparison to the phononic crystals (PnCs) Cupertino et al. (2024); Beccari et al. (2022); Ghadimi (2018), the fundamental perimeter modes of polygon structures Bereyhi et al. (2022b) can reach a high-quality factor of $3.6\times 10^{9}$ with their tenfold reduced size. This compact structure is imperative for practical integration with optical microcavities, since submicron gaps are necessary for strong near-field optomechanical coupling Guo et al. (2019). Additionally, soft-clamping of PnCs is restricted to higher order modes. This will limit the applications on practical sensing, which favors the fundamental mode due to its lower stiffness and cleaner spectral background. Importantly, soft-clamped perimeter modes in polygon resonators demonstrate exceptional frequency scalability, with a quality factor $Q$ greater than $10^{8}$, maintained across more than four octaves ($170~{}\text{kHz}-2.5~{}\text{MHz}$). Moreover, the $Q$s of the perimeter modes exhibit a high thermal-noise-limited force sensitivity of $420~{}\text{zN}\sqrt{\text{Hz}}$ at room temperature, which is comparable to that of atomic force microscopy cantilevers operating at millikelvin temperature ($190~{}\text{zN}\sqrt{\text{Hz}}$).
@@
Perimeter modes are advantageous from the perspectives of engineering design and sensing applications. Nevertheless, the current model simulation, which can be matched with experiments and applications, remains relatively preliminary Bereyhi et al. (2022b). Consequently, we develop a comprehensive analytical framework that incorporates the tether torsion mode into the previous model analysis, and take into account the torsional resonance in the tethers. In the simulations and analysis, the polygon resonators are composed of a commercial material, silicon nitride (Si3N4). The detailed stress analysis is applied to the Timoshenko-Gere (TG) theory. The model displays remarkable predictive accuracy, with frequency estimations within 0.1% of FEM (finite element method) results across the entire operational range. Furthermore, our findings indicate that boundary losses are substantially amplified by the coupling between the polygon side bending modes and the tether torsional modes. Importantly, these losses can be tuned through adjusting the tether length, providing a flexible strategy for optimizing quality factors.
This letter is structured as follows. First, we start with two main sources of the dissipation loss in the polygon systems under the narrow-beam approximation. We then derive the vibration of polygon sides, and the supporting tethers, respectively. Consequently, the analytical dissipation dilution can be derived through the modified TG equations by the coupling of these two vibrations, Next, we calculate the $Q$’s of a sample Si3N4 nanomechanical resonator through FEM and compare the results with our analytical results. We can accurately predict the resonant frequencies and the dissipation dilution factors of perimeter modes.
@@
Dissipation dilution can be fundamentally understood through the interaction between a harmonic oscillator and a lossless potential González (2000). The quality factor of a mechanical mode experiencing dissipation dilution can be enhanced due to
@@
Here, $D_{Q}$ is the dissipation dilution factor and $Q_{\text{int}}$ is the material’s intrinsic quality factor inversely proportional to the material loss angle ($Q_{\text{int}}=1/\phi$) Saulson (1990). The dilution factor $D_{Q}$ of the $n^{\text{th}}$ perimeter mode adopts the general dissipation dilution form of the flexural modes of a stressed mechanical resonatorUnterreithmeier et al. (2009):
@@
with total stored energy $\langle W_{\text{total}}\rangle$, lossy part $\langle W_{\text{lossy}}\rangle$. $\alpha_{n}$ and $\beta_{n}$ are coefficients determined by the resonator’s boundary curvature and the distributed bending over the rest of the mode, respectively. The strain parameter $\lambda=\frac{h}{l}\sqrt{\frac{E}{12\sigma}}$ depends on the resonator length $l$, thickness $h$, Young’s modulus $E$, and stress $\sigma$.
To study the soft-clamped modes in polygon resonators, we start with a four-interconnected beam structure as a general example, as illustrated in Fig. 1. Each beam is supported by two tethers, located at the vertices of the polygon resonator. According to the symmetry of the whole system, the displacement profile of the fundamental perimeter mode is illustrated with a degenerate symmetric vibration pattern in Fig. 2. The transverse reaction forces on opposing tethers demonstrate phase cancelation, resulting in negligible net bending moments. However, the out-of-plane tangential displacement components generate a non-zero torque about the tether axes, inducing torsional dissipation.
In particular, as a result of the practical design considerations, the narrow-beam approximation is employed with $l/w\gg 1$ for all constituent beams with length $l$ and width $w$ Bereyhi et al. (2022b). This justifies the zero boundary loss term ($\alpha_{n}=0$) in the dissipation kernel. Therefore, the dominant dissipation pathways in the polygon system originate from: (i) bending in the polygonal sides, and (ii) shear strain-induced dissipation in the supporting tethers. To analytically estimate the dissipation dilution factor of perimeter modes, we first derive the equations for the transverse displacement $u(x)$ of the polygon sides, and the torsion angle $\tau$ of the supporting tethers, respectively. Then we take their coupling into account.
@@
To model the bending-dominated dissipation, we analyze the transverse displacement field $u(x)$ along each side of the $N$-sided polygon. Within the Euler-Bernoulli beam framework, $u(x)$ follows the flexural string equation in the high tension limit Schmid et al. (2016):
$-\sigma_{s}(x)\frac{d^{2}u}{dx^{2}}=\rho\omega_{n}^{2}u.$ (3)
where $x$ is the axial coordinate along the polygon sides, $\sigma_{s}$ the stress, $\rho$ the material density, and $\omega_{n}$ the degenerate symmetric frequency of the $n^{\text{th}}$ flexural mode. The transverse displacement $u(x)$ can be well approximated by a sinusoidal profile $u(x)=u_{0}\sin{(n\pi x/l_{s})}$ with polygon side length $l_{s}$. The degenerate symmetric frequency $\omega_{n}$ can be determined by substituting $u(x)$ into Eq. 3:
@@
The torsion angle $\tau$ can be determined by the modified Timoshenko-Gere equation incorporating a stress term (see the supplementary material for the derivation of the modified TG equation):
@@
where $\sigma_{t}=\sigma_{xx}$ denotes the axial stress along the tether, $x,y,z$ are the cartesian coordinates, $I_{p}=\iint_{A}(y^{2}+z^{2})\mathrm{d}A$ is the geometric moment of inertia, $I_{\psi}=\iint_{A}\psi^{2}\mathrm{d}A$ ($\psi$ $\approx-yz$ the warping function Chopin and Filho (2019)), $A$ the cross section with thickness $h$ and tether width $w_{t}$, $C$ the torsional rigidity ($C=\iint_{A}G[(\frac{\partial\psi}{\partial y}-z)^{2}+(\frac{\partial\psi}{\partial z}+y)^{2}]\mathrm{d}A$), and $G$ the shear modulus. We found that both Eq. 5 and the Euler-Bernoulli beam equation share the same structure and mode shape representation (Schmid et al. (2016) [p46], and Eq. S5). Consequently, the approximation method, used for high-stress conditions in Eq. 3, is equally applicable to Eq. 5, which can reduce to
$-(C+\sigma_{t}I_{y})\frac{d^{2}\tau}{dx^{2}}=\rho I_{\text{p}}\omega_{n}^{2}\tau,$ (6)
where the axial stress $\sigma_{t}$ contributes to the enhanced effective torsional rigidity of the tether through the term $C+\sigma_{t}I_{y}$. To solve Eq. 6, we assume a trial solution $\tau(x)=\tau_{0}\sin{(kx+\phi_{0})}$ with the boundary condition $\tau(0)=0$. The wave number $k$ is found to be
$k=\sqrt{I_{\text{p}}\sigma_{s}\frac{\pi^{2}}{l_{s}^{2}}\frac{1}{C+\sigma_{t}I_{y}}}.$ (7)
@@
$u^{\prime}(0)=\sin[\frac{(N-2)\pi}{2N}]\tau(x)=\cos{(\frac{\pi}{N})}\tau_{0}\sin{(kl_{t})}.$ (8)
Here, $l_{t}$ is the tether length. Eq. 8 demonstrates that the modal gradient at branch nodes $u^{\prime}(0)$ is determined by both the cosine of the polygon side number $N$ and $\sin{(kl_{t})}$. Since the system’s tensile energy scales quadratically with $u^{\prime}(0)$, theoretical $Q$ maximization requires: (i) increasing the number of polygon sides, and (ii) achieving $\sin{(kl_{t})}=1$. However, the former has been shown to offer a limited practical improvement in dissipation dilution Bereyhi et al. (2022a); Fedorov et al. (2020). Importantly, the latter provides a novel insight: when all material parameters except $l_{t}$ are fixed (wave number $k=\text{const}$), the modal gradient $u^{\prime}(0)$ and consequently the dissipation dilution factor $D_{Q}$ can be tuned by adjusting $l_{t}$.
@@
[…]
where $w_{s}$ is the polygon side width. The lossy energy comprises two components. The distributed bending energy along the polygon sides is
[…]
@@
[…]
Substituting Eq. 9-11 into Eq. 2, and defining $\lambda_{l}=\frac{h}{l_{s}}\sqrt{\frac{E}{12\sigma_{s}}}$ yield a dissipation dilution factor of
@@
The first term in Eq. 12 is due to the distributed bending across the polygon sides, and sets a higher bound on the dissipation dilution factor. The second term depends on $\lambda_{l}^{2}$ and is related to the torsional deformation of the supporting tethers. Notably, the term $\sin^{2}{(kl_{t})}$ in Eq. 14 implies that the dissipation dilution factor $D_{Q,n}$ has a lower bound 0 when $kl_{t}$ is an integer multiple of $\pi$. In this case, the polygon sides exhibit zero transverse displacement, resulting in a system where the total energy manifests exclusively as dissipative torsional energy. To achieve the optimal dilution factor, $l_{t}$ should be chosen to ensure $kl_{t}=\pi/2$.
@@
As an illustrative case, we consider a polygon resonator characterized by the following geometric parameters: vertex count $N=4$, side length $l_{s}=1400~{}\upmu\mathrm{m}$, polygon side width $w_{s}=300~{}\mathrm{nm}$, support length $l_{t}=700~{}\upmu\mathrm{m}$, and support width $w_{t}=400~{}\mathrm{nm}$. The resonator structure is made of a $100~{}\mathrm{nm}$ commercial silicon nitride thin film with Young’s modulus $E=250~{}\mathrm{GPa}$ , mass density $\rho=3100~{}\mathrm{kg/m^{3}}$, Poisson’s ratio $\nu=0.23$, and deposition stress $\sigma_{0}=1~{}\mathrm{GPa}$.
@@
To demonstrate the precision of the theory model, we presented in Fig. 3 a comparative analysis of resonant frequencies and dissipation dilution factors for 20 modes of the polygon resonator, obtained through both the finite element method and analytical modeling. The results demonstrate that the analytical model achieves exceptional accuracy in predicting resonant frequencies, with relative errors below 0.1%. For most modes, the theoretically estimated dilution factors show quantitatively good agreement with FEM results. However, anomalies are observed in the 7th and 14th perimeter modes due to internal resonance effects, where energy transfers from primary flexural modes to coupled torsional modes (comparison between Insets 1, 3 and 4 reveals distinct torsional features). This leads to significant $Q$-factor degradation in primary modes Asadi et al. (2021). Furthermore, the 11th and 12th mode exhibit deviation from theoretical predictions owing to bending-torsion mode coupling in the tethers. Remarkably, the dilution factor of the fundamental mode shows an approximately two orders of magnitude enhancement compared to that of a simple doubly-clamped beam of identical dimensions.
Figure 4: Dissipation dilution factor $D_{\text{Q}}$ for different tether lengths. blue dots and orange dots in d represent FEM simulation and theory model presented in this paper, respectively. The fundamental, second order, and third order modes are presented in Inset a, b, and c, corresponding to the lowest blue dots from left to right.

### v2

# A Predictive Model for the Dissipation Dilution of Soft-Clamped Modes in Polygon Resonators

@@ Abstract
Polygon resonators are promising candidates for nanomechanical applications due to their compact architecture and high force sensitivity. Here, we develop an analytical framework to predict the resonance frequencies and dissipation dilution factors $D_{Q}$ of polygon perimeter modes by extending the Timoshenko-Gere equation to incorporate the tensile stress. The model identifies two dominant dissipation mechanisms: distributed bending in the polygon sides and torsional deformation in the supporting tethers. We reveal that dissipation dilution in these resonators scales as $1/\lambda^{2}$, distinct from the conventional $1/\lambda$ dependence associated with boundary bending loss. Furthermore, we demonstrate that the torsional loss can be suppressed by tailoring the torsion angle of the supporting tethers. The analytical predictions are validated by finite element simulations, providing a predictive framework for designing high-$Q$ polygon resonators for cavity optomechanics and precision sensing.
@@
Recent advancements have facilitated ultra-high quality factor ($Q$) nanomechanical resonators, attributed to the phenomenon known as ”dissipation dilution” Tsaturyan (2017); González (2000). This phenomenon originates from the interplay between geometric nonlinearity of strain in deformations and static strain Fedorov et al. (2019). Notably, dissipation dilution significantly minimizes energy loss in the mechanical system through geometric engineering and stress enhancement, rather than relying solely on material optimization Tsaturyan (2017); Fedorov et al. (2019). This opens up exciting avenues for device design.
@@
Investigating the structural design and forces inherent in nanomechanical systems has revealed that the mode curvature at boundaries is a major source of the mechanical energy loss Yu et al. (2012); Tsaturyan (2017). Reducing the amplitude of the mechanical modes at the clamps is an effective way to minimize the energy loss in these systems. Several innovative techniques, referred to as ”soft clamping” Tsaturyan (2017), have been developed specifically for the clamping points to achieve the low-energy-loss. Previous studies have concentrated on tapered clamping Bereyhi et al. (2019), hierarchical structuring Bereyhi et al. (2022b); Fedorov et al. (2020), phononic bandgap engineering Cupertino et al. (2024); Beccari et al. (2022); Ghadimi (2018), and perimeter modes in polygon resonators Bereyhi et al. (2022a).
@@
Remarkably, the perimeter modes in polygon resonators have emerged as promising candidates for integrated quantum optomechanical systems Bereyhi et al. (2022a). In the tapered-clamping approach Bereyhi et al. (2019), enhancement of $Q$ is challenging, as it is constrained to the material’s yield stress-to-deposition stress ratio. In the hierarchical structuring approach Bereyhi et al. (2022b); Fedorov et al. (2020), the torsional losses in cascaded branchings hinder the resonators from attaining the ”clampless” limit. Furthermore, in comparison to the phononic crystals (PnCs) Beccari et al. (2022); Cupertino et al. (2024); Ghadimi (2018), the fundamental perimeter modes of polygon structures Bereyhi et al. (2022a) can reach a high-quality factor of $3.6\times 10^{9}$ with their tenfold reduced size.This compact architecture is advantageous for practical integration with optical microcavities, as it facilitates the maintenance of stable submicron gaps, a prerequisite for achieving strong near-field optomechanical coupling Guo et al. (2019). Additionally, soft clamping in PnCs is typically realized in defect-localized bandgap modes. While bandgap engineering can provide excellent spectral isolation, practical finite-size devices may still host multiple localized defect modes and, depending on boundaries and fabrication disorder, occasional spurious resonances within or near the nominal bandgap. Importantly, soft-clamped perimeter modes in polygon resonators demonstrate exceptional frequency scalability, with a quality factor $Q$ greater than $10^{8}$, maintained across more than four octaves ($170~\text{kHz}-2.5~\text{MHz}$). Moreover, the $Q$s of the perimeter modes exhibit a high thermal-noise-limited force sensitivity of $420~\text{zN}\sqrt{\text{Hz}}$ at room temperature, which is comparable to that of atomic force microscopy cantilevers operating at millikelvin temperature ($190~\text{zN}\sqrt{\text{Hz}}$). Furthermore, the co-integration of perimeter-mode resonators with on-chip optical cavities suppresses both thermomechanical and readout noises, enabling the systematic study of intrinsic frequency fluctuations. The resulting nonlinear optomechanical transduction yields signals with significantly reduced frequency noise, providing a robust platform for precise frequency tracking Kharbanda et al. (2025).
Perimeter modes are advantageous from the perspectives of engineering design and sensing applications. Nevertheless, the current model simulation, which can be matched with experiments and applications, remains relatively preliminary Bereyhi et al. (2022a). Consequently, we develop a comprehensive analytical framework that incorporates tether torsion modes into the existing models and takes into account the torsional resonance in the tethers. In the simulations and analysis, the polygon resonators are composed of a commercial material, silicon nitride (Si3N4). The detailed stress analysis is applied to the Timoshenko-Gere (TG) theory. The model displays remarkable predictive accuracy, with frequency estimations within 0.1% of FEM (finite element method) results across the entire operational range. Furthermore, our findings indicate that boundary losses are substantially amplified by the coupling between the polygon side bending modes and the tether torsional modes. Importantly, these losses can be tailored by adjusting the tether length, providing a flexible strategy for optimizing quality factors.
@@
This article is structured as follows. First, we identify two dominant dissipation mechanisms in polygon systems under the narrow-beam approximation. We then derive the vibration of polygon sides, and the supporting tethers, respectively. Consequently, the analytical dissipation dilution can be derived through the modified TG equations by the coupling of these two vibrations. Next, we calculate the $Q$’s of a sample Si3N4 nanomechanical resonator through FEM and compare the results with our analytical results. We can accurately predict the resonance frequencies and the dissipation dilution factors of perimeter modes.
@@
Dissipation dilution can be fundamentally understood through the interaction between a harmonic oscillator and a lossless potential González (2000). The quality factor of a mechanical mode subject to dissipation dilution is given by
@@
Here, $D_{Q}$ is the dissipation dilution factor, and $Q_{\text{int}}$ is the material’s intrinsic quality factor, which is inversely proportional to the material loss angle ($Q_{\text{int}}=1/\phi$) Saulson (1990). For the $n^{\text{th}}$ resonator mode, the dissipation dilution factor $D_{Q}$ can be written in the general form as Unterreithmeier et al. (2009); Fedorov et al. (2019, 2020):
@@
where $\langle W_{\text{total}}\rangle$ denotes the total stored energy, and $\langle W_{\text{lossy}}\rangle$ is the lossy part. The coefficients $\alpha_{n}$ and $\beta_{n}$ are determined by the resonator’s boundary curvature and the distributed bending over the rest of the mode, respectively. The strain parameter $\lambda=\frac{h}{l}\sqrt{\frac{E}{12\sigma}}$ depends on the resonator length $l$, thickness $h$, Young’s modulus $E$, and tensile stress $\sigma$.
To study soft-clamped modes in polygon resonators, we start with a four-interconnected beam structure as a general example, as illustrated in Fig. 1. Each beam is supported by two tethers, located at the vertices of the polygon. Due to the underlying symmetry, the fundamental perimeter mode possesses a symmetric displacement profile, as shown in Fig. 2. The transverse reaction forces on opposing tethers demonstrate phase cancellation, resulting in negligible net bending moments. However, the out-of-plane tangential displacement components generate a non-zero torque about the tether axes, inducing torsional dissipation.
@@
In particular, as a result of practical design considerations, the narrow-beam approximation is employed with $l/w\gg 1$ for all constituent beams with length $l$ and width $w$ Bereyhi et al. (2022a). Moreover, previous work has shown that the ratio of the torsional energy stored in a beam to its tensile energy scales as $\lambda^{2}$ Fedorov et al. (2020). This justifies setting the boundary-loss coefficient to be zero ($\alpha_{n}=0$) in the dissipation kernel. Therefore, the dominant dissipation pathways in polygon resonators arise from: (i) bending of the polygonal sides and (ii) shear strain-induced dissipation in the supporting tethers. To analytically estimate the dissipation dilution factor of the perimeter modes, we first derive the equations for the transverse displacement $u(x)$ of the polygon sides and the torsion angle $\tau$ of the supporting tethers, respectively. We then incorporate the coupling between these modes into the model.
@@
To model the bending-dominated dissipation, we analyze the transverse displacement field $u(x)$ along each side of the $N$-sided polygon. Within the Euler-Bernoulli beam framework, $u(x)$ satisfies the flexural string equation in the high tension limit Schmid et al. (2016):
$\sigma_{s}(x)\frac{d^{2}u}{dx^{2}}=\rho\omega_{n}^{2}u.$ (3)
where $x$ is the axial coordinate along the polygon sides, $\sigma_{s}$ the stress, $\rho$ the material density, and $\omega_{n}$ the degenerate symmetric frequency of the $n^{\text{th}}$ flexural mode. The stress distribution $\sigma_{s}$ is determined by force balance and the constancy of the path-length between the two clamped vertices of the $N$-sided polygon Bereyhi et al. (2022a):
$\sigma_{s}=\frac{\sigma_{0}(1-\nu)[1+2r_{l}\sin{(\pi/N)}]}{1+4(r_{l}/r_{w})\sin^{2}{}(\pi/N)},$ (4)
with $\sigma_{0}$ representing the initial deposition stress, $\nu$ the Poisson’s ratio, $r_{l}=l_{s}/l_{t}$ the ratio of the polygon side length to the tether length, and $r_{w}=w_{s}/w_{t}$ the ratio of the polygon side width to the tether width. The assumption of perfectly sinusoidal deformation holds under the narrow beam condition ($l\gg w$), reflecting the kinematic constraints of ”soft-clamped” boundaries at vertices. The transverse displacement $u(x)$ can be well approximated by a sinusoidal profile $u(x)=u_{0}\sin{(n\pi x/l_{s})}$ with polygon side length $l_{s}$. The degenerate symmetric frequency $\omega_{n}$ can be determined by substituting $u(x)$ into Eq. 3:
@@
The torsion angle $\tau$ is determined by the modified TG equation, incorporating the axial stress term (see Appendix A for the full derivation of the modified TG equation):
@@
where $\sigma_{t}=\sigma_{xx}$ denotes the axial stress along the tether, $x,y,z$ are the cartesian coordinates, $I_{p}=\iint_{A}(y^{2}+z^{2})\mathrm{d}A$ is the geometric moment of inertia, $I_{\psi}=\iint_{A}\psi^{2}\mathrm{d}A$ ($\psi$ $\approx-yz$ the warping function Chopin and Filho (2019)), $A$ the cross section with thickness $h$ and tether width $w_{t}$, $C$ the torsional rigidity ($C=\iint_{A}G[(\frac{\partial\psi}{\partial y}-z)^{2}+(\frac{\partial\psi}{\partial z}+y)^{2}]\mathrm{d}A$), and $G$ the shear modulus. We note that Eq. 6 is structurally analogous to the Euler-Bernoulli beam equation in both the structure and mode shape representation (Schmid et al. (2016) (p46), and Eq. S5). Consequently, the approximation method, used for high-stress conditions in Eq. 3, is equally applicable to Eq. 6, which can be simplified to
$(C+\sigma_{t}I_{y})\frac{d^{2}\tau}{dx^{2}}=\rho I_{\text{p}}\omega_{n}^{2}\tau,$ (7)
where the axial stress $\sigma_{t}$ represents the enhanced effective torsional rigidity of the tether through the term $C+\sigma_{t}I_{y}$. To solve Eq. 7, we assume a trial solution $\tau(x)=\tau_{0}\sin{(kx+\phi_{0})}$ with the boundary condition $\tau(0)=0$, therefore the wave number $k$ is identified as
$k=\sqrt{I_{\text{p}}\sigma_{s}\frac{n^{2}\pi^{2}}{l_{s}^{2}}\frac{1}{C+\sigma_{t}I_{y}}}.$ (8)
@@
$u^{\prime}(0)=\sin[\frac{(N-2)\pi}{2N}]\tau(l_{t})=\cos{(\frac{\pi}{N})}\tau_{0}\sin{(kl_{t})}.$ (9)
Eq. 9 demonstrates that the modal gradient at branch nodes, $u^{\prime}(0)$, is determined by both the cosine of the polygon side number $N$ and $\sin{(kl_{t})}$. Since the system’s tensile energy scales quadratically with $u^{\prime}(0)$, theoretical $Q$ maximization requires: (i) increasing the number of polygon sides, and (ii) achieving $\sin{(kl_{t})}=1$. However, the former has been shown to offer a limited practical improvement in dissipation dilution Bereyhi et al. (2022b); Fedorov et al. (2020). Importantly, the latter provides a novel insight: when all material parameters except $l_{t}$ are fixed (wave number $k=\text{const}$), the modal gradient $u^{\prime}(0)$ - and consequently the dissipation dilution factor $D_{Q}$ - can be optimized by adjusting $l_{t}$.
@@
$\langle W_{\text{tens}}\rangle=\frac{N}{2}\sigma_{s}w_{s}h\int_{0}^{l_{s}}[u^{\prime}(x)]^{2}dx.$ (10)
The lossy energy comprises two components. The distributed bending energy along the polygon sides is
$\langle W_{\text{bend}}\rangle=\frac{N}{2}\frac{Ew_{s}h^{3}}{12}\int_{0}^{l_{s}}[u^{{}^{\prime\prime}}(x)]^{2}dx,$ (11)
@@
$\langle W_{\text{tors}}\rangle=N\frac{Gw_{t}h^{3}}{12}\int_{0}^{l_{t}}(\tau^{\prime}(x))^{2}dx.$ (12)
Substituting Eqs. 10 - 12 into Eq. 2 and defining $\lambda_{l}=\frac{h}{l_{s}}\sqrt{\frac{E}{12\sigma_{s}}}$ yields a dissipation dilution factor of
@@
The first term in Eq. 13 is due to the distributed bending across the polygon sides, and sets an upper bound on the dissipation dilution factor. The second term is related to the torsional deformation of the supporting tethers. Notably, the term $\sin^{2}{(kl_{t})}$ in Eq. 15 implies that the dissipation dilution factor $D_{Q,n}$ has a lower bound 0 when $kl_{t}$ is an integer multiple of $\pi$. In this case, the polygon sides exhibit zero transverse displacement, resulting in a system where the total energy manifests exclusively as dissipative torsional energy. To achieve the optimal dilution factor, $l_{t}$ should be chosen to ensure $kl_{t}=\pi/2$. It is crucial that the torsional loss part $D_{Q,\text{tors}}$ depends inversely on $\lambda_{l}^{2}$, thereby establishing a scaling law fundamentally different from the $1/\lambda$ dependence of clamping loss in conventional dissipation dilution Fedorov et al. (2020); Bereyhi et al. (2022b, 2019). This insight promises a new paradigm for engineering ultra-high-Q resonators, one that circumvents the need for the large dimensions historically required by phononic crystals to open a bandgap and approach the ”clampless” limit Ghadimi (2018).
@@
As an illustrative case, we consider a polygon resonator characterized by the following geometric parameters: vertex count $N=4$, side length $l_{s}=1400~\upmu\mathrm{m}$, polygon side width $w_{s}=300~\mathrm{nm}$, support length $l_{t}=700~\upmu\mathrm{m}$, and support width $w_{t}=400~\mathrm{nm}$. The resonator structure is made of a $100~\mathrm{nm}$ commercial Si3N4 thin film with Young’s modulus $E=250~\mathrm{GPa}$, mass density $\rho=3100~\mathrm{kg/m^{3}}$, Poisson’s ratio $\nu=0.23$, and deposition stress $\sigma_{0}=1~\mathrm{GPa}$.
Figure 4: Dissipation dilution factors of fundamental perimeter modes with $N=4$. (a) $D_{Q}$ dependence on the tether length, with $l_{s}=700~\upmu\mathrm{m}$, $w_{s}=300~\mathrm{nm}$, and $w_{t}=400~\mathrm{nm}$. The orange triangles and light blue circles represent the FEM simulation results and the analytical predictions, respectively. (b) Torsion angle profiles for the tether at lengths of 3920, 7890, and $11810~\upmu\mathrm{m}$. The results from the FEM simulation are shown in orange and are compared against sinusoidal references (light blue). (c) $D_{Q}$ dependence on the polygon width, with $l_{s}=700~\upmu\mathrm{m}$, $l_{t}=700~\upmu\mathrm{m}$, and $w_{t}=400~\mathrm{nm}$. (d) Distribution of $D_{Q}$ for $l_{s}=700~\upmu\mathrm{m}$ and $w_{s}=300~\mathrm{nm}$. (e) Distribution of $D_{Q}$ for $l_{t}=700~\upmu\mathrm{m}$ and $w_{t}=400~\mathrm{nm}$.
[…]
@@
To demonstrate the precision of the analytical model, we present in Fig. 3 a comparative analysis of resonance frequencies and dissipation dilution factors for 20 modes of the polygon resonator, as obtained through both the FEM and analytical modeling. The analytical model predicts the resonance frequencies with relative errors less than 0.1%. For most modes, the analytically estimated dilution factors are in good quantitative agreement with the FEM results. However, anomalies are observed in the $7^{\mathrm{th}}$ and $14^{\mathrm{th}}$ perimeter modes due to internal resonance effects, where energy transfers from primary flexural modes to coupled torsional modes (comparison between Insets a, b, and d reveals distinct torsional features). This energy transfer leads to significant $Q$-factor degradation in primary modes Asadi et al. (2021). In addition, the $11^{\mathrm{th}}$ and $12^{\mathrm{th}}$ modes (inset c) exhibit deviation from theoretical predictions due to the coupling of the bending-torsion mode in the tethers. Remarkably, the dilution factor of the fundamental mode is approximately two orders of magnitude higher than that of a simple doubly-clamped beam whose dimensions match those of a single polygon side.